I need some help with a college algebra problem I don't understand,
It says to solve by the elimination method
0.3x-0.2y=4
0.2x+0.5y=-21/23
It says the solution is an ordered pair and to use integers or fractions for any numbers in the expression, can someone please help me and Thank You

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

solve by the elimination method
0.3x-0.2y=4
0.2x+0.5y=-21/23
-----------------------
Multiply thru each equation by 1000 to get:
300x - 200y = 4000
200x + 500y = -913
-------------------------
Let's say you want to eliminate the x-terms:
Multiply thru the 1st equation by 2
Multiply thru the 2nd equation by 3
600x - 400y = 8000
600x +1500y = -2739
--------------------
Subtract the 1st from the 2nd to get:
1900y = -10,739
y = -5.65
----------------
Substitute into 300x - 200y = 4000 to solve for "x":
300x - 200*-5.65 = 4000
300x = 2870
x = 9.57
-------------------
These problems are so messy because they assume you
are using a calculator to do the arithmetic.
============================================
Cheers,
Stan H.

Answer by richwmiller(17219)

0.3x-0.2y=4
0.2x+0.5y=-21/23
multiply the first equation by 5
.15x-y=20
and the second by 2
.4x+y=-42/23
.15x-y=20
add the two equations
.19x=(460-42)/23
.19x=418/23
x=(418/23)/.19
x=95.6522
0.3x-0.2y=4
0.2x+0.5y=-21/23
multiply the first by 2
0.6x-0.4y=8
and the second by 3
0.2x+0.5y=-21/23
0.6x+1.5y=-63/23
0.6x-0.4y=8
subtract
1.9y=-8-63/23
y=(247/23)*(1.9)
y=5.65217